OpenStudy (anonymous):

2/3 divided by (-1/4)

OpenStudy (lgbasallote):

\[\huge \frac{\frac 23}{\frac{-1}{4}} \implies \frac 23 \times \frac{-4}{1}\] does that help?

OpenStudy (anonymous):

yes!!

OpenStudy (lgbasallote):

wonderful!

OpenStudy (anonymous):

you are